OpenStudy (samigupta8):

The plane x+2y+z=6 cuts x,y and z axes at A,B and C.Find x coordinate of circumcentre of ∆ABC

OpenStudy (priyar):

you can find the three coordinates.. then use that to find the circumcentre

OpenStudy (samigupta8):

Really! do you know the formula for that? For circumcentre part

OpenStudy (priyar):

no i don't know any formula but there are two methods..

OpenStudy (samigupta8):

One method i definitely tried.Bt ended up nowhere.

OpenStudy (priyar):

which one ? intersection of perpendicular bisectors?

OpenStudy (samigupta8):

The way in which we equate the radius of the circle knowing the three vertices and a centre assumed to be (a,b,c)

OpenStudy (priyar):

oh okay, but the shorter method is you can find angle A and the solve

OpenStudy (samigupta8):

Other way is finding or tho center but i don't know that too.

OpenStudy (priyar):

there is one formula that involves sin2A.. have you heard of it?

OpenStudy (priyar):

but the perpendicular bisectors intersection will surely work

OpenStudy (samigupta8):

We can't apply it here ,i think so (sin2A formula) becoz this is a 3-D figure.

OpenStudy (samigupta8):

I m just now getting how can you apply perpendicular bisectors also becoz we never know the equation of lines here. Do we?

OpenStudy (samigupta8):

I m just now getting how can you apply perpendicular bisectors also becoz we never know the equation of lines here. Do we?

Parth (parthkohli):

\[(6,0,0), (0, 3, 0), (0, 0, 6)\]By symmetry, the \(x\) and \(z\) coordinates of the center will be the same. So assume the center to be \((a, b, a)\).\[(a-6)^2 + b^2 + a^2 = a^2 + (b-6)^2 + a^2 \]\[\Rightarrow -6(2a-6) = 6(2b-6)\]\[\Rightarrow 3 - a = b-3\]\[\Rightarrow a + b = 6\]\[(a-6)^2 + (a-6)^2+a^2 = a^2+(a-3)^2 + a^2 \]Finally, \(a = 7/2\) and \(b = 5/2\). The circumcenter is \((7/2, 5/2, 7/2)\)

OpenStudy (priyar):

we can find the eqs of line using the coordinates

OpenStudy (priyar):

If you want to know this method (per. bis.) here is a link: https://www.youtube.com/watch?v=kwooTzS01PA

OpenStudy (samigupta8):

How you did this? a^2+(b-6)^2+a^2 in the first equation I got this a^2+(b-3)^2+a^2

OpenStudy (samigupta8):

@parthkohli i literally don't know how did you came up with values of a and b I just got an equation which is 9=4a-2b

OpenStudy (priyar):

Even i didn't understand.. @parthkohli can explain from the symmetry part?

OpenStudy (samigupta8):

Did you get the same equation as i wrote above. Further as far as i see,he has done a typo in that.

OpenStudy (priyar):

Even i didn't understand.. @parthkohli can explain from the symmetry part?